# Dimension-independent convergence rate of propagation of chaos and numerical analysis for McKean-Vlasov stochastic differential equations with coefficients nonlinearly dependent on measure

### v1

@@ 3.1 Strong quantitative propagation of chaos
${\rm d}X_{t}^{i,N}=$ $a(X_{t}^{i,N}){\rm d}t+A\left(\frac{1}{N}\sum_{j=1}^{N}\kappa(X_{t}^{i,N},X_{t}^{i,N})\right){\rm d}t+B\left(\frac{1}{N}\sum_{j=1}^{N}\zeta(X_{t}^{i,N},X_{t}^{i,N})\right){\rm d}W^{i}_{t},$ (10)
which is the corresponding interacting particle system of (5)-(6), and the following quantitive convergence rate of PoC follows as an immediate by-product of Theorem 2.1
@@
Let $p\geq 2$, it holds under Assumptions 2.1-2.3 that
$\mathbb{E}\left(\sup_{0\leq t\leq T}|x_{t}^{i}-X_{t}^{i,N}|^{p}\right)\leq CN^{-p/2}.$
According to the definition of $\mathbb{W}_{p}$-metric, we can get the following result directly.
Corollary 3.2.
For $p\geq 2$, under the conditions of Theorem 3.1,
$\sup_{1\leq i\leq N}\mathbb{W}_{p}(\mu_{t}^{x^{i}},\mu_{t}^{X^{i,N}})\leq CN^{-1/2},$
where $\mu_{t}^{X^{i,N}}$ denotes the marginal law of $X^{i,N}$ at the time $t$.
@@
[38] showed that $\mathbb{W}_{2}(\mu_{t}^{x^{i}},\mu_{t}^{X^{i,N}})=\mathcal{O}(N^{-1/2})$. While in the case of Lipschitz continuous dependence in measure in the $\mathbb{W}_{2}$-metric, the rate of strong PoC deteriorates with the dimension $d$, for instance, the PoC rate under the $L^{2}$-norm presented in [14, 12, 10] is
@@
the rate under the $L^{p}$-norm for $p>0$ and $q>p$ given in [18, 27] is
@@
It can be observed that the PoC rate presented above depends on the dimension of the system and decays at approximately $O(N^{-1/2p})$ under the $\mathbb{W}_{p}$-distance when $p>d/2$. Note that, in this work, we treat a case of MV-SDEs with coefficients Lipschitz continuous dependence in measure, and prove a dimension-independent PoC rate of $\mathcal{O}(N^{-1/2})$ under the $\mathbb{W}_{p}$-distance. This paper significantly contributes to this area by extending previous findings.
3.2 Strong convergence rate of the tamed EM method
@@
in (7), and establish the convergence of the tamed EM method for the particle system (5). It is easy to known that
@@
${\rm d}X_{t}^{i,N,\Delta}=$ $\frac{a(X_{\delta_{t}}^{i,N,\Delta})}{1+\sqrt{\Delta}|a(X_{\delta_{t}}^{i,N,\Delta})|}{\rm d}t+A\left(\frac{1}{N}\sum_{j=1}^{N}\kappa(X_{\delta_{t}}^{i,N,\Delta},X_{\delta_{t}}^{i,N,\Delta})\right){\rm d}t+B\left(\frac{1}{N}\sum_{j=1}^{N}\zeta(X_{\delta_{t}}^{i,N,\Delta},X_{\delta_{t}}^{i,N,\Delta})\right){\rm d}W^{i}_{t},$ (13)
which is in accordance with the continuous version of the tamed EM method for (5).
@@
Let $p\geq 2$, under Assumptions 2.1-2.3, it holds that
@@
For any $i\in\{1,\dots,N\}$, using the Itô formula to $|X_{t}^{i,N,\Delta}|^{p}$ with $X_{t}^{i,N,\Delta}$ satisfies (13), we have
@@
Then for any $T_{1}\in[0,T]$, since $X^{i,N,\Delta},i=1,2,\dots,N$ are identically distributed, according to the Young and B-D-G inequalities, we have
@@
$+\mathbb{E}\int_{0}^{T_{1}}\left|B\left(\frac{1}{N}\sum_{j=1}^{N}\zeta(X_{\delta_{t}}^{i,N,\Delta},X_{\delta_{t}}^{j,N,\Delta})\right)(W_{t}^{i}-W_{\delta_{t}}^{i})\right|^{p/2}\Delta^{-p/4}{\rm d}t$
@@
Substituting this inequality into (3.2), it can be derived that
@@
Moreover, since $|\tilde{a}(x)|\leq|a(x)|$, using Remark 2.1, one has
@@
$\sup_{1\leq i\leq N}\mathbb{E}\left(\sup_{0\leq t\leq T}|a(X_{t}^{i,N,\Delta})-\tilde{a}(X_{t}^{i,N,\Delta})|^{p}\right)\leq C\Delta^{p/2}.$
@@
$\leq$ $\mathbb{E}\left(\sup_{0\leq t\leq T}|a(X_{t}^{i,N,\Delta})|^{p}\right)\Delta^{p/2}\leq C\Delta^{p/2}\left(1+\mathbb{E}\sup_{0\leq t\leq T}|X_{t}^{i,N,\Delta}|^{p(q+1)}\right)\leq C\Delta^{p/2}.$
@@
$\sup_{1\leq i\leq N}\mathbb{E}\left(\sup_{0\leq t\leq T}|x_{t}^{i}-X_{t}^{i,N,\Delta}|^{p}\right)\leq CN^{-p/2}+C\Delta^{p/2}.$
@@ 4 Numerical simulation
for a fixed time step size $\Delta=2^{-10}$, other parameters will be specified in the following. By performing a linear regression on the logarithm of the error and the number of particles using the least squares method, we give the images of the order of convergence.
@@
Consider the scalar nonlinear MV-SDE (5) with coefficients
$a(x)=x-x^{3},\quad\kappa(x,y)=y,\quad\zeta(x,y)=x,\quad A(x)=\frac{1}{1+{\rm e}^{-x}},\quad B(x)=x.$
@@
$x_{t}=x_{0}+\int_{0}^{t}(x_{s}-x_{s}^{3}){\rm d}s+\int_{0}^{t}\frac{1}{1+{\rm e}^{-\mathbb{E}(x_{s})}}{\rm d}s+\int_{0}^{t}\mathbb{E}(x_{s}){\rm d}W_{s}.$
And it is easy to verify that $a(x),A(x),B(x),\kappa(x,y),\zeta(x,y)$ satisfy Assumptions 2.1-2.4. The corresponding interacting particle system is
$X_{t}^{i,N}=$ $x_{0}+\int_{0}^{t}(X_{s}^{i,N}-(X_{s}^{i,N})^{3}){\rm d}s+\int_{0}^{t}\left(1+\exp\left\{-\frac{1}{N}\sum_{j=1}^{N}X_{s}^{j,N}\right\}\right)^{-1}{\rm d}s+\frac{1}{N}\sum_{j=1}^{N}\int_{0}^{t}X_{s}^{j,N}{\rm d}W_{s}^{i},$
@@
$X_{n+1}^{i,N}=$ $X_{n}^{i,N}+\frac{X_{n}^{i,N}-(X_{n}^{i,N})^{3}}{1+\sqrt{\Delta}|X_{n}^{i,N}-(X_{n}^{i,N})^{3}|}\Delta+\left(1+\exp\left\{-\frac{1}{N}\sum_{j=1}^{N}X_{n}^{i,N}\right\}\right)^{-1}\Delta+\frac{1}{N}\sum_{j=1}^{N}X_{n}^{j,N}\Delta W_{n}^{i}.$
The orders of convergence with respect to the number of particles $N$ are shown in Figure 1. The proxy solution takes $N=2^{12}$. The numerical solutions take $\bar{N}=2^{8},2^{9},2^{10},2^{11}$.
Figure 1: convergence rate of the PoC
@@
In this example, we test the $L^{p}$ rate of PoC in different dimensions for $p=2,4,6$ and compare the findings to the theoretical upper bounds established by (11). For system (5)-(6), we make the following choices: Let $d\geq 2$, $x=(x_{1},...,x_{d})\in\mathbb{R}^{d}$, $(W_{t})_{0\leq t\leq T}$ is a 1-dim Brownian motion, the initial condition is a vector distributed according to $d$-independent $N(0,1)$-random variables, the coefficients are taken to be $a(x)=x-x^{3}=(x_{1}-x_{1}^{3},\dots,x_{d}-x_{d}^{3})^{\rm T}$,
$\kappa(x,y)=y,\quad\zeta(x,y)=$ $\left(\begin{array}[]{c}\sin(x_{1})+x_{2}+\cdots+x_{d}+y_{1}\\ x_{1}+\sin(x_{2})+\cdots+x_{d}+y_{2}\\ \cdots\\ x_{1}+x_{2}+\cdots+\sin(x_{d})+y_{d}\end{array}\right),$
and $A(x)=\sin(x)=(\sin(x_{1}),\dots,\sin(x_{d})^{\rm T},B(x)=\cos(x)=(\cos(x_{1}),\dots,\cos(x_{d}))^{\rm T}$. We performed PoC rate tests in the sense of $L^{2}$, $L^{4}$ and $L^{6}$, respectively. When $p=2$, we simulated the equations for dimensions $d=2$ (then $p>d/2$), $d=4(p=d/2)$ and $d=6(p\in[2,d/2))$, when $p=4$, we simulated the equations for dimensions $d=4$ ($p>d/2$), $d=8(p=d/2)$ and $d=10(p\in[2,d/2))$, and when $p=6$, we simulated the equations for dimensions $d=4,12,14$. It can be found in Figures 2, 3 and 4 that the PoC rates observed numerically are all around $\mathcal{O}(N^{-1/2})$ and do not depend on the dimension of the solution.
@@
Figure 2: $L^{2}$ convergence rate of the PoC.

### v2

@@ 4 Dimension-independent rate of propagation of chaos
${\rm d}X_{t}^{i,N}=$ $a(X_{t}^{i,N}){\rm d}t+f(X_{t}^{i,N},\hat{\mu}_{t}^{X,N}){\rm d}t+g(X_{t}^{i,N},\hat{\mu}_{t}^{X,N}){\rm d}W^{i}_{t},$
which is the corresponding interacting particle system of (4), and the following quantitive convergence rate of PoC follows as an immediate by-product of Theorem 3.1.
@@
Let $p\geq 2$, it holds under Assumptions 3.1-3.3 that
$\mathbb{E}\left(\sup_{0\leq t\leq T}|x_{t}^{i}-X_{t}^{i,N}|^{p}\right)\leq CN^{-\frac{p}{2}},$
where $x_{t}^{i}$ and $X_{t}^{i,N}$ are the solutions of MV-SDEs (4) with coefficient given by (6) and its corresponding interacting particle system, respectively.
In alignment with Theorems 3.2 and 3.3, when the coefficients in equation (4) are specified by (7)-(8) or (9)-(10), the same convergence rate of PoC can be achieved. To avoid redundancy, we omit the explicit statement of the corresponding theorem here.
@@
[38] showed that $\mathbb{E}\left(\sup_{0\leq t\leq T}|x_{t}^{i}-X_{t}^{i,N}|\right)\leq CN^{-\frac{1}{2}}$. While in the case of the coefficients of MV-SDEs are Lipschitz continuous in measure, the rate of strong PoC deteriorates with the dimension $d$, for instance, the PoC rate under the $L^{2}$-norm presented in [10, 12, 14] is
@@
and the rate under the $L^{p}$-norm for $p>0$ and $q>p$ given in [18, 27] is
@@
It can be observed that the PoC rate presented above depends on the dimension of the system and decays at approximately $O(N^{-1/2p})$ when $p>d/2$. At this point, if the coefficients simultaneously satisfy some stronger regularity conditions, the rate $1/2$ was previously established in [13, Lemma 5.1] and [39, Theorem 4.2]. However, in this work, we treat some special cases of nonlinear MV-SDEs with coefficients do not claim strong regularity, this is obviously a different set of sufficient conditions compared to the existing results.
5 Strong convergence rate of the time-stepping method
The convergence of the time-stepping method has been well studied in many literatures, such as tamed EM method ([14]), truncated EM method ([18]), adaptive EM method ([35]), and so on. The convergence rate of the time-stepping scheme is not the concern of this paper, but in order to give a computable numerical scheme for MV-SDEs, here we introduce the tamed EM method and utilize it in the next section to compute the solution of the test equations, which then verifies the convergence order of the PoC. Here we only give the numerical scheme of equation (4) when the coefficients are taken to be (6) and verify Assumption 3.3. Finally, the order of convergence of the numerical method is obtained according to Theorem 3.1.
@@ 5 Strong convergence rate of the time-stepping method
in (5), and establish the convergence of the tamed EM method for the particle system (4) with coefficients given by (6). It is easy to known that
@@
${\rm d}X_{t}^{i,N,\Delta}=$ $\frac{a(X_{\delta_{t}}^{i,N,\Delta})}{1+\sqrt{\Delta}|a(X_{\delta_{t}}^{i,N,\Delta})|}{\rm d}t+A\left(\frac{1}{N}\sum_{j=1}^{N}\kappa(X_{\delta_{t}}^{i,N,\Delta},X_{\delta_{t}}^{j,N,\Delta})\right){\rm d}t+B\left(\frac{1}{N}\sum_{j=1}^{N}\zeta(X_{\delta_{t}}^{i,N,\Delta},X_{\delta_{t}}^{j,N,\Delta})\right){\rm d}W^{i}_{t},$ (28)
which is in accordance with the continuous version of the tamed EM method for (4) with coefficients given by (6).
@@
Let $p\geq 2$, under Assumptions 3.1-3.3, it holds that
@@
For any $i\in\{1,\dots,N\}$ and $p>2$, using the Itô formula to $|X_{t}^{i,N,\Delta}|^{p}$ with $X_{t}^{i,N,\Delta}$ satisfies (28), we have
@@
Then for any $T_{1}\in[0,T]$, since $\{X_{t}^{i,N,\Delta}\}_{1\leq i\leq N}$ are identically distributed, according to the Young and B-D-G inequalities, we have
@@
$+\mathbb{E}\int_{0}^{T_{1}}\left|B\left(\frac{1}{N}\sum_{j=1}^{N}\zeta(X_{\delta_{t}}^{i,N,\Delta},X_{\delta_{t}}^{j,N,\Delta})\right)(W_{t}^{i}-W_{\delta_{t}}^{i})\right|^{\frac{p}{2}}\Delta^{-p/4}{\rm d}t$
@@
Substituting this inequality into (5), it can be derived that
@@
Similarly, we can also get the 2nd moment boundedness of the numerical solution. Moreover, since $|\tilde{a}(x)|\leq|a(x)|$, using Remark 3.1, one has
@@
$\sup_{1\leq i\leq N}\mathbb{E}\left(\sup_{0\leq t\leq T}|a(X_{t}^{i,N,\Delta})-\tilde{a}(X_{t}^{i,N,\Delta})|^{p}\right)\leq C\Delta^{\frac{p}{2}}.$
@@
$\leq$ $\mathbb{E}\left(\sup_{0\leq t\leq T}|a(X_{t}^{i,N,\Delta})|^{p}\right)\Delta^{\frac{p}{2}}\leq C\Delta^{\frac{p}{2}}\left(1+\mathbb{E}\sup_{0\leq t\leq T}|X_{t}^{i,N,\Delta}|^{p(q+1)}\right)\leq C\Delta^{\frac{p}{2}}.$
@@
$\sup_{1\leq i\leq N}\mathbb{E}\left(\sup_{0\leq t\leq T}|x_{t}^{i}-X_{t}^{i,N,\Delta}|^{p}\right)\leq CN^{-\frac{p}{2}}+C\Delta^{\frac{p}{2}}.$
@@ 6 Numerical simulation
for a fixed time step size $\Delta=2^{-10}$, other parameters will be specified in the following. By performing a linear regression on the logarithm of the error and the number of particles using the least squares method, we give the images of the order of convergence. In this section, let $\phi(x)$ denote $\left(\phi(x_{1}),\phi(x_{2}),\dots,\phi(x_{d})\right)^{\rm T}$ for $x\in\mathbb{R}^{d}$ and let $\phi(x)=\left(\phi(x_{ij})\right)_{d\times d}$ for $x\in\mathbb{R}^{d\times d}$.
@@
First, we consider a scalar nonlinear MV-SDE (4) with coefficients taking in the form (6) under
$a(x)=x-x^{3},\quad A(x)=\frac{1}{1+{\rm e}^{-x}},\quad\kappa(x,y)=\arctan(x+y),\quad B(x)=\sin(x),\quad\zeta(x,y)=\sqrt{x^{2}+y^{2}},$
@@
$x_{t}=x_{0}+\int_{0}^{t}(x_{s}-x_{s}^{3}){\rm d}s+\int_{0}^{t}\frac{1}{1+{\rm exp}\{-\mathbb{E}[\arctan(z+x_{s})]\big{|}_{z=x_{s}}\}}{\rm d}s+\int_{0}^{t}\sin\left(\mathbb{E}[\sqrt{z^{2}+x_{s}^{2}}]\big{|}_{z=x_{s}}\right){\rm d}W_{s}.$
It is easy to verify that $a(x),A(x),B(x),\kappa(x,y),\zeta(x,y)$ satisfy Assumptions 3.1-3.3. The corresponding interacting particle system is
$X_{t}^{i,N}=$ $x_{0}+\int_{0}^{t}(X_{s}^{i,N}-(X_{s}^{i,N})^{3}){\rm d}s+\int_{0}^{t}\left(1+\exp\left\{-\frac{1}{N}\sum_{j=1}^{N}\arctan(X_{s}^{i,N}+X_{s}^{j,N})\right\}\right)^{-1}{\rm d}s$
$+\int_{0}^{t}\sin\left(\frac{1}{N}\sum_{j=1}^{N}\sqrt{(X_{s}^{i,N})^{2}+(X_{s}^{j,N})^{2}}\right){\rm d}W_{s}^{i},$
@@
$X_{t_{n+1}}^{i,N}=$ $X_{t_{n}}^{i,N}+\frac{X_{t_{n}}^{i,N}-(X_{t_{n}}^{i,N})^{3}}{1+\sqrt{\Delta}|X_{t_{n}}^{i,N}-(X_{t_{n}}^{i,N})^{3}|}\Delta+\left(1+\exp\left\{-\frac{1}{N}\sum_{j=1}^{N}\arctan(X_{t_{n}}^{i,N}+X_{t_{n}}^{j,N})\right\}\right)^{-1}\Delta$
$+\sin\left(\frac{1}{N}\sum_{j=1}^{N}\sqrt{(X_{t_{n}}^{i,N})^{2}+(X_{t_{n}}^{j,N})^{2}}\right)\Delta W_{n}^{i},$
where $t_{n}=n\Delta,n=0,1,\dots.$ The orders of convergence with respect to the number of particles $N$ are shown in Figure 1. The proxy solution takes $N=2^{11}$. The numerical solutions take $\bar{N}=2^{7},2^{8},2^{9},2^{10}$.
Figure 1: $L^{p}$ convergence rate of PoC for Example 1.
@@
In this example, we test the $L^{p}$ rate of PoC in different dimensions for $p=2,4$ and compare the findings to the theoretical upper bounds established by (26). For system (4) with coefficients taking the form of (6), we make the following choices: Let $d\geq 2$, $x=(x_{1},...,x_{d})\in\mathbb{R}^{d}$, $(W_{t})_{0\leq t\leq T}$ is a d-dim Brownian motion, the initial condition is a vector distributed according to $d$-independent $N(0,1)$-random variables, the coefficients are taken to be $a(x)=x-x^{3}:=(x_{1}-x_{1}^{3},\dots,x_{d}-x_{d}^{3})^{\rm T}$,
$\kappa(x,y)=\left(\begin{array}[]{c}{\rm sign}(x_{1})|x_{1}+y_{1}|\\ {\rm sign}(x_{2})|x_{2}+y_{2}|\\ \vdots\\ {\rm sign}(x_{d})|x_{d}+y_{d}|\end{array}\right),\quad\zeta(x,y)=\left(\begin{array}[]{cccc}\sqrt{x_{1}^{2}+y_{1}^{2}}&x_{2}&\cdots&x_{d}\\ x_{1}&\sqrt{x_{2}^{2}+y_{2}^{2}}&\cdots&x_{d}\\ \vdots&\vdots&&\vdots\\ x_{1}&x_{2}&\cdots&\sqrt{x_{d}^{2}+y_{d}^{2}}\end{array}\right),$
and $A(x)=\sin(x),B(x)=\cos(x)$. The tamed EM scheme is given by
$X_{t_{n+1}}^{i,N}=$ $X_{t_{n}}^{i,N}+\frac{(X_{t_{n}}^{i,N}-(X_{t_{n}}^{i,N})^{3})\Delta}{1+\sqrt{\Delta}|X_{t_{n}}^{i,N}-(X_{t_{n}}^{i,N})^{3}|}+\sin\left(\frac{1}{N}\sum_{j=1}^{N}\kappa(X_{t_{n}}^{i,N},X_{t_{n}}^{j,N})\right)\Delta+\cos\left(\frac{1}{N}\sum_{j=1}^{N}\zeta(X_{t_{n}}^{i,N},X_{t_{n}}^{j,N})\right)\Delta W_{n}^{i}.$
We performed PoC rate tests in the sense of $L^{2}$, $L^{4}$ and $L^{6}$, respectively. When $p=2$, we simulated the equations for dimensions $d=2$ (then $p>d/2$), $d=4(p=d/2)$ and $d=6(p\in[2,d/2))$, when $p=4$, we simulated the equations for dimensions $d=4$ ($p>d/2$), $d=8(p=d/2)$ and $d=10(p\in[2,d/2))$, and when $p=6$, we simulated the equations for dimensions $d=4,12,14$. It can be found in Figures 2 and 3 that the PoC rates observed numerically are all around $\mathcal{O}(N^{-1/2})$ and do not depend on the dimension of the solution.
@@
Figure 2: $L^{2}$ convergence rate of PoC for Example 2.